### v2

# Extremely closed subgroups and a variant on Glauberman’s $Z^{*}$-theorem

Abstract.
Let $G$ be a finite group and let $H$ be a subgroup of $G$. We say that $H$ is extremely closed in $G$ if $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G.$ In this paper, we determine the structure of finite groups with an extremely closed abelian $p$-subgroup for some prime $p$. In particular, we show that if $G$ contains such a subgroup $H$, then $G=N_{G}(H)O_{p^{\prime}}(G).$ This is a variant on the celebrated Glauberman’s $Z^{*}$-Theorem.
Key words and phrases:
extremely closed, strongly closed, weakly closed
2020 Mathematics Subject Classification:
Primary 20D05, 20D15
1. Introduction
It is an important problem in finite group theory to determine whether a finite group is simple or not. Many non-simplicity criteria have been obtained in the literature. Among those is the celebrated Glauberman’s $Z^{*}$-theorem. To state this theorem, we need some definitions. Let $G$ be a finite group and let $p$ be a prime. Let $x\in G$ be a $p$-element and let $P$ be a Sylow $p$-subgroup of $G$ containing $x$. We say that $x$ is isolated in $P$ with respect to $G$ if $x^{G}\cap P=\{x\}$, that is, $x$ is not conjugate in $G$ to any element in $P$ except for $x$ itself. Here, $x^{G}$ denotes the conjugacy class of $G$ containing $x$. We say that $x$ is isolated in $G$ if $x$ is isolated in some Sylow $p$-subgroup of $G$ containing it. Glauberman’s $Z^{*}$-theorem [18] states that if $x\in G$ is an isolated involution in $G$, then $G=C_{G}(x)O_{2^{\prime}}(G).$ The proof of this theorem depends on the modular representation theory and is independent of the classification of finite groups. Recall that for a prime $p$, $O_{p^{\prime}}(G)$ is the largest normal $p^{\prime}$-subgroup of $G$. Extending this fundamental theorem to all primes, Glauberman’s $Z_{p}^{*}$-theorem states that if $x\in G$ is an isolated $p$-element, then $G=C_{G}(x)O_{p^{\prime}}(G)$. For various proofs of this theorem, see [1, 19, 28, 40]. Note that all of these proofs depend on the classification of finite simple groups. For many equivalent statements of this theorem, see [30]. Also, see [14] for some variant of Glauberman’s $Z_{p}^{*}$-theorem.
In this paper, we introduce the so-called extremely closed subgroup and obtain some new factorization of finite groups similar to Glauberman’s $Z^{*}_{p}$-theorem which gives some non-simplicity criteria for finite groups. Let $G$ be a finite group and let $H\leqslant M$ be subgroups of $G$. We say that $H$ is extremely closed in M with respect to $G$ if $\langle H,H^{g}\rangle\cap M=H$ for all $g\in G-M,$ and that $H$ is extremely closed in $G$ if $H$ is extremely closed in $N_{G}(H)$ with respect to $G$. Trivially, if $H$ is a normal or self-normalizing subgroup of $G,$ then $H$ is extremely closed in $G.$
If $G$ is a finite group, we write $Z(G)$ for the center of $G$ and $\Phi(G)$ for the Frattini subgroup of $G$, that is, the intersection of all maximal subgroups of $G$. Furthermore, if $H$ is a subgroup of $G$, then $\langle H^{G}\rangle$ is the normal closure of $H$ in $G$. We next compare our definition of extremely closed subgroups with other known embedding properties of subgroups of finite groups.
The first motivation for our definition comes from work of Flavell [9] on the generation of finite groups with maximal subgroups of maximal subgroups. In particular, a triple $(G,M,H)$ with $H\leqslant M\leqslant G$ is called a $\gamma$-triple if $H<M<G$ and $\langle H,g\rangle\cap M=H$ for all $g\in G-M.$ If $H$ is maximal in $M$ and $M$ is maximal in $G$ and moreover, $G$ cannot be generated by any two conjugates of $H$, then $(G,M,H)$ is a $\gamma$-triple. Clearly, if $(G,M,H)$ is a $\gamma$-triple, then $H$ is extremely closed in $M$ with respect to $G$. The converse is not true by the example below.
Example 1.
Let $G={\mathrm{S}}_{4}$ be the symmetric group of degree $4$. Let $H=\langle(1,2,3)\rangle$ and $M=N_{G}(H)\cong{\mathrm{S}}_{3}$. Then $\langle H^{G}\rangle={\mathrm{A}}_{4}$ and $\langle H^{G}\rangle\cap M=H.$ So $\langle H,H^{g}\rangle\cap M=H$ for all $g\in G$, and hence $H$ is extremely closed in $M$ with respect to $G$. However, let $g=(1,3,2,4)\in G-M$. Then $\langle H,g\rangle=G$ and so $\langle H,g\rangle\cap M=M\neq H$. Therefore $(G,M,H)$ is not a $\gamma$-triple.
For the second motivation, following Hawkes and Humphreys, a subgroup $M$ of a finite group $G$ is said to have property CR (character restriction) if every irreducible complex character of $M$ is the restriction of a character of $G$. In [31], the authors studied finite solvable groups with a CR subgroup and the general cases were considered by Isaacs in [32]. One important property of a CR-subgroup $M$ of a finite group $G$ is that if $H\unlhd M$, then $\langle H^{G}\rangle\cap M=H$ (See [32, Proposition 1.1]). Berkovich [3] called a triple $(G,M,H)$ with $H\unlhd M\leqslant G$ special in $G$ if $\langle H^{G}\rangle\cap M=H.$ (In [36], Li calls $H$ an NE-subgroup of $G$ if $(G,N_{G}(H),H)$ is special in $G$.) In [32], Isaacs showed that if $P$ is a Sylow $p$-subgroup of $G$, where $p$ is a prime and assume that $N_{G}(P)$ satisfies CR in $G$, then $N_{G}(P)$ has a normal complement in $G$. This result was extended by Berkovich [3], where he showed that if both triples $(G,N_{G}(P),P)$ and $(G,N_{G}(P),\Phi(P))$ are special in $G$, then $N_{G}(P)$ has a normal complement in $G$. This gives a character theory free proof of Isaacs’ result mentioned earlier. Observe that if a triple $(G,N_{G}(H),H)$ is special in $G,$ then $H$ is extremely closed in $G$. However, the converse is not true.
Example 2.
Let $G=P:H$ be a semidirect product of $H$ and $P,$ where $H=\langle a\rangle$ is a cyclic group of order $2$ and $P\cong 3_{+}^{1+2}$ is an extraspecial group of order $27$ with exponent $3,$ so
$P=\langle x,y,z|z=[x,y],x^{3}=y^{3}=z^{3}=1=[x,z]=[y,z]\rangle$
and $H$ acts on $P$ via $x^{a}=x^{-1},y^{a}=y^{-1}$ and $z^{a}=z.$ Then $N_{G}(H)=C_{G}(H)=H\langle z\rangle,$ and $\langle H^{G}\rangle=G.$ For every $g\in G-N_{G}(H),$ we have $T=\langle H,H^{g}\rangle$ is a dihedral group of order $6,$ so $N_{T}(H)=H,$ and hence $H$ is extremely closed in $G$ but $(G,N_{G}(H),H)$ is not a special triple since $\langle H^{G}\rangle\cap N_{G}(H)=H\langle z\rangle\neq H.$
Finally, we mention the last inspiration for our new embedding property. Let $H\leqslant M$ be subgroups of a finite group $G$. Recall that $H$ is said to be strongly closed in $M$ with respect to $G$ if, whenever $a^{g}\in M,$ where $a\in H,g\in G$, then $a^{g}\in H.$ This is equivalent to saying that $M\cap H^{g}\leqslant H$ for all $g\in G.$ Furthermore, we say that $H$ is strongly closed in $G$ if $H$ is strongly closed in $N_{G}(H)$ with respect to $G$. Noting that in [4], $H$ is called an $\mathscr{H}$-subgroup of $G$ if $H$ is strongly closed in $G$. If $H=\langle x\rangle$ is cyclic of order $2$, then $H$ is strongly closed in $G$ if and only if $x$ is isolated in $G$. Finite groups with a strongly closed $p$-subgroup are determined in [15, 16, 20]. It is easy to see that if $H$ is extremely closed in $G$, then $H$ is strongly closed in $G$.
Example 3.
Let $G=\mathrm{U}_{3}(4)$. By [20], if $P$ is a Sylow $2$-subgroup of $G$, then $H=Z(P)=\Phi(P)$ is a strongly closed abelian $2$-subgroup of $G$. Using GAP [17], we can find $g\in G$ of order $15$ such that $\langle H,H^{g}\rangle\cong{\mathrm{A}}_{5}$, $N_{G}(H)\cong P\langle g\rangle$ and $\langle H,H^{g}\rangle\cap N_{G}(H)\cong{\mathrm{A}}_{4}\neq H$. Thus $H$ is not extremely closed in $G$.
Finally, we recall the definition of weakly closed subgroups. Let $H\leqslant M$ be subgroups of a finite group $G$. We say that $H$ is weakly closed in $M$ with respect to $G$ if, whenever $H^{g}\subseteq M,$ where $g\in G,$ then $H^{g}=H.$ It is easy to see that if $H$ is strongly closed in $M$ with respect to $G$, then $H$ is weakly closed in $M$ with respect to $G$. Moreover, when $H$ is cyclic of prime order, these two concepts coincide.
We know that if $H\unlhd M\leqslant G$ and $\langle H^{G}\rangle\cap M=H$, then $H$ is extremely closed in $M$ with respect to $G$. In our first result, we show that in certain cases, the converse holds.
Theorem 1.
Let $G$ be a finite group and let $H\leqslant M$ be subgroups of $G$. Suppose that $\langle H,H^{g}\rangle\cap M=H$ for all $g\in G.$ If $H$ is maximal in $M$ and $M$ is maximal in $G,$ then $\langle H^{G}\rangle\cap M=H.$
Let $H\leqslant M$ be subgroups of a finite group $G$. Recall that $H$ is called a weak second maximal subgroup of $G$ if there exists a maximal subgroup $M$ of $G$ such that $H$ is maximal in $M$. Moreover, $H$ is called a second maximal subgroup of $G$ if $H\neq G$ and it is maximal in every maximal subgroup of $G$ containing it. In [9], Flavell shows that if $G$ is a finite nonabelian simple group and $H$ is a weak second maximal subgroup of $G$, then $G=\langle H,g\rangle$ for some $g\in G.$ And in [10], it is shown that if $H$ is a second maximal subgroup of a finite nonabelian simple group $G$, then $G=\langle H,H^{g}\rangle$ for some $g\in G.$ Using the same argument as in the proof of Theorem 1, we obtain the following generation result for finite nonabelian simple groups.
Theorem 2.
Let $G$ be a finite nonabelian simple group. Let $M$ be a maximal subgroup of $G$ and let $H$ be a normal subgroup of $M$ of prime index. Then $G=\langle H,H^{g}\rangle$ for some $g\in G.$
[…]
Let $G$ be a finite group and let $p$ be a prime. We now focus on extremely closed $p$-subgroups. Let $H$ be an extremely closed $p$-subgroup of $G$. We first assume that $H=\langle x\rangle$ is cyclic of order $p.$ If $p=2,$ then it is not hard to see that $x$ is isolated in $G$ or equivalently $H$ is strongly closed in $G$ and so $G=C_{G}(x)O_{2^{\prime}}(G)$ by Glauberman’s $Z^{*}$-Theorem. In particular, $G$ is not simple. When $p$ is odd, we note that there exists a simple group with a weakly closed or strongly closed subgroup of order $p,$ for instance, when a Sylow $p$-subgroup of $G$ is cyclic. However, when $H$ is extremely closed in $G$, the subgroups generated by any two distinct conjugates of $H$ are Frobenius groups. By applying a result due to B. Fischer [8] on the Frobenius automorphisms, we can show that $(G,N_{G}(H),H)$ is special in $G,$ and then we obtain a factorization similar to that of Glauberman’s $Z_{p}^{*}$-Theorem.
Theorem 3.
Let $G$ be a finite group and let $p$ be an odd prime divisor of the order of $G$. Let $H$ be a cyclic subgroup of order $p$ of $G$. If $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G$, then $\langle H^{G}\rangle\cap N_{G}(H)=H$ and $G=N_{G}(H)O_{p^{\prime}}(G).$
It would be nice to have a proof of Theorem 3 which does not rely on the classification of finite simple groups. When $H$ is an abelian $p$-group, again $p$ is odd, we also obtain the same results as in Theorem 3. In the proof, we make use of a result due to Guest [25] on the characterization of solvable radical of finite groups and the classification of finite groups with strongly closed subgroups by Flores and Foote [15].
Theorem 4.
Let $G$ be a finite group and let $p$ be an odd prime. Let $H$ be an abelian $p$-subgroup of $G$. If $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G,$ then $\langle H^{G}\rangle\cap N_{G}(H)=H$ and $G=N_{G}(H)O_{p^{\prime}}(G).$
Example 2 above shows that the first conclusion of Theorem 4 does not hold for $p=2$. However, the second claim remains true.
Theorem 5.
Let $G$ be a finite group and let $H$ be an abelian $2$-subgroup of $G$. Assume that $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G.$ Then $G=N_{G}(H)O_{2^{\prime}}(G).$
Note that the proof of Theorem 5 depends on the classification of finite groups with an abelian strongly closed subgroup by Goldschmidt [20] which is independent of the classification of finite simple groups. Remark that we cannot drop the hypothesis that $H$ is abelian, since the simple group $G=\mathrm{L}_{2}(17)$ has a self-normalizing Sylow $2$-subgroup $P$ which is nonabelian and so $P$ is clearly an extremely closed $2$-subgroup of $G$. Note that if $|H|=2$, then Theorem 5 is just Glauberman’s $Z^{*}$- theorem
Recall that for a finite group $G$, the solvable radical of $G$, denoted by $R(G)$, is the largest normal solvable subgroup of $G$. As a consequence of Theorems 4 and 5, we obtain the following.
Corollary 1.
Let $G$ be a finite group and let $p$ be a prime. If $H$ is an extremely closed abelian $p$-subgroup of $G$, then $H\subseteq R(G)$.
By an application of Burnside’s normal $p$-complement theorem and the solvability of finite groups admitting a fixed point free coprime group action, if $H$ satisfies the hypothesis of the corollary, then $\langle H,H^{g}\rangle$ is solvable for all $g\in G.$ Thus if $x\in H$, then $\langle x,x^{g}\rangle$ is solvable for all $g\in G.$ By the main results in [24, 25], if $p\geq 5$, then $x\in R(G)$ and hence $H\leqslant R(G)$. Thus the above corollary only provides new result when $p=2$ or $3$.
In general, if $x$ is a $p$-element and $\langle x,x^{g}\rangle$ is $p$-solvable for all $g\in G$, then it is not true that $x\in R_{p}(G)$, where $R_{p}(G)$ is the $p$-solvable radical of $G$, that is, $R_{p}(G)$ is the largest normal $p$-solvable subgroup of $G$. For a counterexample, consider $G=\mathrm{U}_{3}(3)$ and $x\in G$ a transvection, so $x$ has order $3$ and the conjugacy class of $G$ containing $x$ has size $56$, then we can check that $\langle x,x^{g}\rangle$ is isomorphic to either $\langle x\rangle$ or ${\mathrm{SL}}_{2}(3)$ for every $g\in G.$ Clearly $\langle x,x^{g}\rangle$ is $3$-nilpotent and hence it is $3$-solvable for every $g\in G.$ There is also a counterexample when $p=2$ since if $x\in G$ is an involution, then $\langle x,x^{g}\rangle$ is $2$-nilpotent for every $g\in G.$ Recall that a finite group $G$ is $p$-nilpotent if it has a normal $p$-complement for some prime $p.$ On the other hand, it is proved in [12] that if $P$ is a Sylow $p$-subgroup of a finite group $G$ for some prime $p,$ then $G$ is $p$-solvable if and only if $\langle P,g\rangle$ is $p$-solvable for all $g\in G$. Generalizing this result, we can prove the following.
Theorem 6.
Let $G$ be a finite group and let $p$ be a prime. Let $P$ be a Sylow $p$-subgroup of $G$. Then $G$ is $p$-solvable if and only if $\langle P,P^{g}\rangle$ is $p$-solvable for all $g\in G.$
Our notation is standard. For finite group theory, we follow [21] and [35] and for finite simple groups, we follow the notation in [34].
For the organization of the paper, we collect some useful results in Section 2. We will prove Theorems 1-5 and the corollary in Section 3 and the last theorem will be proved in Section 4.
2. Preliminaries
Let $G$ be a finite group. Recall that the Fitting subgroup of $G$, denoted by $F(G)$, is the largest nilpotent normal subgroup of $G$. The layer of $G$, denoted by $E(G)$, is the product of all components of $G$, where a component of $G$ is a subnormal quasisimple subgroup of $G$. A finite group $L$ is quasisimple if $L$ is perfect and $L/Z(L)$ is a nonabelian simple group. The generalized Fitting subgroup of $G$, denoted by $F^{*}(G)$, is defined by $F^{*}(G)=F(G)E(G).$ As usual, if $H\leqslant G$, then $N_{G}(H)$ and $C_{G}(H)$ denote the normalizer and centralizer of $H$ in $G$, respectively. Finally, a finite group $G$ is almost simple with socle $S$ if there exists a finite nonabelian simple group $S$ such that $S\unlhd G\leqslant\operatorname{Aut}(S).$
Recall that a subgroup $H$ of $G$ is called pronormal (resp. abnormal) in $G$ if for any $g\in G,$ $H^{g}=H^{u}$ for some $u\in\langle H,H^{g}\rangle,$ (resp. $\langle H,H^{g}\rangle=\langle H,g\rangle$). The first lemma is obvious, for completeness, we will include a proof here.
Lemma 2.1.
Let $G$ be a finite group. Let $H$ be a pronormal subgroup of $G$ and let $N\unlhd G.$ Then the following hold.
$(i)$ If $N\unlhd G$ and $P\in{\mathrm{Syl}}_{p}(N)$, then $P$ is pronormal in $G$.
$(ii)$ $N_{G}(H)$ is abnormal in $G$.
$(iii)$ If $H\leqslant N\unlhd G$, then $G=N_{G}(H)N$.
$(iv)$ If $H\leqslant L\leqslant G$, then $H$ is pronormal in $L$.
$(v)$ If $H$ is subnormal in $K$, where $K\leqslant G,$ then $H\unlhd K$.
$(vi)$ If $L=\langle H^{G}\rangle$, then $\langle H^{L}\rangle=L.$
Proof.
$(i)$ Let $g\in G$. Sine $P\leqslant N\unlhd G,$ we have $\langle P,P^{g}\rangle\leqslant N.$ As $P\in{\mathrm{Syl}}_{p}(N),$ it follows that $P\in{\mathrm{Syl}}_{p}\langle P,P^{g}\rangle$ and hence by Sylow’s Theorem, $P^{g}=P^{u}$ for some $u\in\langle P,P^{g}\rangle.$ Thus $P$ is pronormal in $G.$
$(ii)$ Assume that $H$ is pronormal in $G.$ Let $M=N_{G}(H)$ and let $g\in G$. By definition, $H^{g}=H^{u},$ for some $u\in\langle H,H^{g}\rangle,$ whence $gu^{-1}\in N_{G}(H)=M.$ Since $u\in\langle H,H^{g}\rangle\leqslant\langle M,M^{g}\rangle,$ we have $g\in\langle M,M^{g}\rangle$ and so $\langle M,M^{g}\rangle=\langle M,g\rangle$. Hence $M$ is abnormal in $G$.
$(iii)$ Let $g\in G.$ We have $H^{g}=H^{u},$ where $u\in\langle H,H^{g}\rangle\leqslant N$ as $H\leqslant N\unlhd G.$ Thus $gu^{-1}\in N_{G}(H)$ and so $g\in N_{G}(H)N.$
$(iv)$ This is obvious.
$(v)$ By $(iv),$ it suffices to show that if $H$ is subnormal and pronormal in $G$ then $H\unlhd G.$ In fact, we only need to prove the following: if $H\unlhd K\unlhd G$ and $H$ is pronormal in $G$ then $H\unlhd G.$ By applying $(iii),$ we have $G=N_{G}(H)K.$ However, as $H\unlhd K,$ $K\leqslant N_{G}(H)$ and so $G=N_{G}(H).$
$(vi)$ Since $H\leqslant L=\langle H^{G}\rangle\unlhd G,$ $G=N_{G}(H)L$ by $(iii)$ and thus
$L=\langle H^{G}\rangle=\langle H^{N_{G}(H)L}\rangle=\langle H^{L}\rangle\leqslant L.$
Therefore $L=\langle H^{L}\rangle.$ This completes the proof of the lemma. ∎
We next deduce some properties of extremely closed subgroups.
Lemma 2.2.
Let $G$ be a finite group and let $H$ be a $p$-subgroup of $G$ for some prime $p$. Let $N\unlhd G$ and assume that $H$ is extremely closed in $G$. Let $\overline{G}=G/N.$ Then the following hold.
$(i)$ $H$ is pronormal in $G.$
$(ii)$ If $H\leqslant L$, then $H$ is extremely closed in $L$.
$(iii)$ $N_{\overline{G}}(\overline{H})=\overline{N_{G}(H)}.$
$(iv)$ If $H$ is abelian, then $\langle H,H^{g}\rangle=HO_{p^{\prime}}\langle H,H^{g}\rangle$, for every $g\in G$.
$(v)$ $\overline{H}$ is extremely closed in $\overline{G}$.
$(vi)$ If $H\leqslant Q\leqslant G,$ where $Q$ is a $p$-group, then $N_{G}(Q)\leqslant N_{G}(H).$ In particular, if $H\leqslant P\in{\mathrm{Syl}}_{p}(G)$, then $P\leqslant N_{G}(H).$
Proof.
$(i)$ Let $g\in G$ and let $T=\langle H,H^{g}\rangle.$ We have $N_{T}(H)=H$ and so $H$ is a Sylow $p$-subgroup of $T.$ As $|H^{g}|=|H|$ and $H^{g}\leqslant T,$ $H^{g}$ is also a Sylow $p$-subgroup of $T$. By Sylow’s theorem, $H^{g}=H^{u}$ for some $u\in T.$ Thus $H$ is pronormal in $G.$
$(ii)$ Let $g\in L$. Then $\langle H,H^{g}\rangle\cap N_{L}(H)=\langle H,H^{g}\rangle\cap N_{G}(H)\cap L=H\cap L=H.$
$(iii)$ It suffices to show that $N_{G}(HN)\leqslant N_{G}(H)N.$ Let $g\in N_{G}(HN).$ Then $H^{g}\leqslant HN$ and hence $T=\langle H,H^{g}\rangle\leqslant HN.$ Since $H$ is pronormal in $G$ by $(i),$ we have $H^{g}=H^{u},$ for some $u\in T\leqslant HN.$ Thus $gu^{-1}\in N_{G}(H)$ whence $g\in N_{G}(H)HN=N_{G}(H)N.$
$(iv)$ Assume that $H$ is abelian. Let $g\in G$ and let $T=\langle H,H^{g}\rangle.$ As $H$ is extremely closed in $G$, $T\cap N_{G}(H)=N_{T}(H)=H.$ Hence $H$ is a self-normalizing abelian Sylow $p$-subgroup of $T.$ The result now follows from Burnside’s normal $p$-complement theorem ([21, Theorem $7.4.3$]).
$(v)$ Applying $(iii),$ we need to show that $\langle H,H^{g}\rangle N\cap N_{G}(H)N=HN$, for all $g\in G.$ Let $T=\langle H,H^{g}\rangle.$ By Dedekind’s Modular law, we have
$TN\cap N_{G}(H)N=N(TN\cap N_{G}(H)).$
Hence, it suffices to show that $TN\cap N_{G}(H)\leqslant HN.$ Let $y=xn\in TN\cap N_{G}(H),$ where $y\in N_{G}(H),x\in T$ and $n\in N.$ We have $H=H^{y}=H^{xn},$ it implies that $H^{x}=H^{n^{-1}}.$ As $N_{T}(H)=H,$ by $(i)$ and Lemma 2.1$(ii),$ $H$ is abnormal in $T,$ and thus
$x\in\langle H,H^{x}\rangle=\langle H,H^{n^{-1}}\rangle\leqslant HN.$
Therefore, $y=xn\in HN.$
$(vi)$ Since $H$ is pronormal and subnormal in $N_{G}(Q),$ by Lemma 2.1$(v),$ $H$ is normal in $N_{G}(Q).$ The remaining claim is obvious. The proof of the lemma is now complete. ∎
Lemma 2.3.
Let $G$ be a finite group, let $N\unlhd G$ and let $H$ be an extremely closed $p$-subgroup of $G$ for some prime $p$. Let $P$ be a Sylow $p$-subgroup of $G$ containing $H$ and let $Q=H\cap N$. Then the following hold.
$(i)$ $H$ is strongly closed in $P$ with respect to $G$.
$(ii)$ $Q$ is strongly closed in $P\cap N$ with respect to $N$.
$(iii)$ If $N\leqslant N_{G}(H),$ then $Q\unlhd G.$
Proof.
First, observe that $R:=P\cap N\in{\mathrm{Syl}}_{p}(N)$ and by Lemma 2.2(vi), $P\leqslant N_{G}(H)$.
$(i)$ For $g\in G$, we have $H^{g}\cap P\leqslant\langle H,H^{g}\rangle\cap N_{G}(H)=H.$ So $H$ is strongly closed in $P$ with respect to $G$.
$(ii)$ For $n\in N,$ we have $Q^{n}\leqslant H^{n}$ and $R\leqslant P\leqslant N_{G}(H)$ and so
$Q^{n}\cap R\leqslant\langle H,H^{n}\rangle\cap N_{G}(H)=H.$
Furthermore, as $Q\leqslant N\unlhd G$ and $R\leqslant N,$ we have $Q^{n}\cap R\leqslant N.$ Hence $Q^{n}\cap R\leqslant H\cap N=Q.$
$(iii)$ Assume that $N\leqslant N_{G}(H).$ For each $g\in G,$ we have
$Q^{g}=Q^{g}\cap N\leqslant\langle H,H^{g}\rangle\cap N_{G}(H)\cap N=H\cap N=Q.$
Hence $Q\unlhd G$ as wanted. The proof is now complete. ∎
We next quote some results that we will need for the proofs of the main theorems.
Lemma 2.4.
Suppose that $G$ is a finite group with $F(G)=1.$ Let $L$ be a component of $G.$ If $x\in G$ such that $x\not\in N_{G}(L)$ and $x^{2}\not\in C_{G}(L)$ then there exists an element $g\in G$ such that $\langle x,x^{g}\rangle$ is not solvable.
Proof.
This is Lemma $1$ in [25]. ∎
Lemma 2.5.
Let $G$ be a finite almost simple group with socle $L.$ Suppose that $x\in G$ is an element of order $p,$ where $p$ is an odd prime. Then one of the following holds:
$(i)$ $\langle x,x^{g}\rangle$ is not solvable for some $g\in G;$
$(ii)$ $p=3$ and $L$ is a finite simple group of Lie type defined over $\mathbb{F}_{3},$ a finite field with $3$ elements, or $L\cong\mathrm{U}_{n}(2),n\geqslant 4.$ Moreover, the Lie rank of $L$ is at least $2$ unless $L\cong\mathrm{U}_{3}(3).$
Proof.
This is Theorem $A^{*}$ in [25]. ∎
Recall that a triple $(G,M,H)$ with $H\unlhd M\leqslant G$ is called a $W$-triple if $M\cap M^{g}\leqslant H$ for all $g\in G-M.$
Lemma 2.6.
Let $G$ be a finite group and $H\unlhd M\leqslant G.$ Then $(G,M,H)$ is a $W$-triple if and only if
$N_{G}(D)\leqslant M\mbox{ for all subgroups $D\leqslant M$ with $D\not\leqslant H.$}$
Proof.
This is Lemma 2.3 in [9]. ∎
Lemma 2.7.
(Wielandt’s Theorem). Let $G$ be a finite group. If $(G,M,H)$ is a $W$-triple, then $G$ contains a normal subgroup $K$ such that $G=MK$ and $M\cap K=H.$ In particular, the triple $(G,M,H)$ is special in $G.$
Proof.
The first claim is in [42] or [38, Exercise 1, p. 347] and the second is in [3, Lemma 9]. ∎
An automorphism $\theta$ of a finite group is Frobenius if each nontrivial power of $\theta$ is fixed point free.
Lemma 2.8.
Let $G$ be a finite group and let $D$ be a conjugacy class of $G$ containing elements of order $>2$. Assume that $G=\langle D\rangle$. Then some element of $D$ induces a Frobenius automorphism on $G^{\prime}$ if and only if each pair of distinct elements in $D$ generates a Frobenius group.
Proof.
This is Satz I in [8]. ∎
We also need the following results.
Lemma 2.9.
Let $G$ be a finite group.
  • $(i)$
    Let $\pi$ be a set of odd primes and suppose that the $\pi$-group $P$ acts as a group of automorphisms on the solvable finite $\pi^{\prime}$-group $G$. Then
    $C_{[G,P]}(P)=\langle C_{[g,P]}(P):g\in G\rangle.$
  • $(ii)$
    Let $\alpha$ be a coprime automorphism of odd order of $G$. Then
    $C_{[G,\langle\alpha\rangle]}(\alpha)=\langle C_{[g,\langle\alpha\rangle]}(\alpha):g\in[G,\langle\alpha\rangle]\rangle.$
Proof.
The first claim is [13, Theorem A] and the second is [14, Theorem 2]. ∎
We will use the next result repeatedly.
Lemma 2.10.
Let $\pi$ be a non-empty set of primes. Let a finite $\pi$-group $Q$ acting fixed point freely on a finite $\pi^{\prime}$-group $R$, that is, $C_{R}(Q)=1$, then $R$ is solvable.
Proof.
For a proof, see Theorem 2.3 in [27]. ∎
We also need the following consequence of Burnside’s normal $p$-complement theorem.
Lemma 2.11.
Let $G$ be a finite group and let $p$ be a prime. Let $H$ be an abelian $p$-subgroup of $G$. Assume that $\langle H^{G}\rangle\cap N_{G}(H)=H$, then $G=N_{G}(H)O_{p^{\prime}}(G).$ Moreover, $N_{G}(H)$ has a normal complement in $G$, which is $O_{p^{\prime}}(\langle H^{G}\rangle)$.
Proof.
Let $H$ be an abelian $p$-subgroup of $G$. Assume that $\langle H^{G}\rangle\cap N_{G}(H)=H$. Let $L=\langle H^{G}\rangle\unlhd G.$ Then $N_{L}(H)=L\cap N_{G}(H)=H$, so $H$ is a self-normalizing abelian Sylow $p$-subgroup of $L$. In particular, $H\leqslant Z(N_{L}(H))$ and hence by Burnside’s normal $p$-complement theorem ([21, Theorem 7.4.3]) $L=HO_{p^{\prime}}(L)$. By Frattini’s argument
$G=N_{G}(H)L=N_{G}(H)O_{p^{\prime}}(L)=N_{G}(H)O_{p^{\prime}}(G).$
The last equality holds since $O_{p^{\prime}}(L)\leqslant O_{p^{\prime}}(G)$.
We now show that $O_{p^{\prime}}(L)$ is a normal complement to $N_{G}(H)$ in $G$. To see this, observe that $L=HO_{p^{\prime}}(L)$ and $H\cap O_{p^{\prime}}(L)=1.$ Note that $O_{p^{\prime}}(L)\unlhd G$. Thus it suffices to show that $N_{G}(H)\cap O_{p^{\prime}}(L)=1.$ Indeed, we have
$N_{G}(H)\cap O_{p^{\prime}}(L)=N_{G}(H)\cap L\cap O_{p^{\prime}}(L)=H\cap O_{p^{\prime}}(L)=1.$
The proof is now complete. ∎
3. Extremely closed abelian $p$-subgroups
We are now ready to prove the main theorems of the paper. We first prove Theorem 1.
Proof of Theorem 1..
[…]
Clearly, if $H\unlhd G$, then the conclusion holds. So, assume that $M=N_{G}(H).$ Suppose that there exists a subgroup $D\leqslant M$ with $D\not\leqslant H$ and $N_{G}(D)\not\leqslant M.$ Take $g\in N_{G}(D)-M.$ Then $D$ normalizes both $H$ and $H^{g}.$ Thus $D\langle H,H^{g}\rangle\leqslant G.$ Since $H$ is normal and maximal in $M$ and $D\not\leqslant H,$ we have $M=HD$ and hence $M\leqslant D\langle H,H^{g}\rangle\leqslant G.$ As $g\not\in N_{G}(H)=M,$ $\langle H,H^{g}\rangle\not\leqslant M$ (otherwise, $H=\langle H,H^{g}\rangle\cap M=\langle H,H^{g}\rangle$). So $G=D\langle H,H^{g}\rangle$ by the maximality of $M.$ We have
$\langle H^{G}\rangle=\langle H^{D\langle H,H^{g}\rangle}\rangle\leqslant\langle H,H^{g}\rangle\leqslant\langle H^{G}\rangle$
and so $\langle H^{G}\rangle=\langle H,H^{g}\rangle.$ Thus $\langle H^{G}\rangle\cap M=\langle H,H^{g}\rangle\cap M=H$ and we are done.
Therefore, we can assume that whenever $D\leqslant M$ with $D\not\leqslant H$ then $N_{G}(D)\leqslant M.$ By Lemma 2.6, $(G,M,H)$ is a $W$-triple and so the result follows from Lemma 2.7. ∎
Proof of Theorem 2.
[…]
We first observe that $H>1$ and $M=N_{G}(H)$. Indeed, if $H=1$, then $M$ is a self-normalizing cyclic subgroup of $G$ of prime order $p$ (since $G$ is nonabelian simple), which is impossible by Burnside’s normal $p$-complement theorem. Therefore, $H>1$ and thus $M=N_{G}(H)$.
We next claim that $(G,M,H)$ is a $W$-triple. Let $D\leqslant M$ with $D\not\leqslant H.$ We will show that $N_{G}(D)\leqslant M$ and thus by Lemma 2.6, $(G,M,H)$ is a $W$-triple as required. Assume by contradiction that $N_{G}(D)\not\leqslant M$ and let $g\in N_{G}(D)-M.$ Then $D$ normalizes both $H$ and $H^{g}$ and thus $D\langle H,H^{g}\rangle$ is a subgroup of $G$ containing $DH=M$. Since $M$ is maximal in $G$, either $D\langle H,H^{g}\rangle=G$ or $D\langle H,H^{g}\rangle=M.$ Assume first that the former case holds. Then $G=\langle H^{G}\rangle=\langle H^{D\langle H,H^{g}\rangle}\rangle\leqslant\langle H,H^{g}\rangle\leqslant G.$ Thus $G=\langle H,H^{g}\rangle$, violating our assumption. Therefore, we must have $M=D\langle H,H^{g}\rangle.$ It follows that $H^{g}\leqslant M.$ Now we have that $M^{g}=(DH)^{g}=DH^{g}\leqslant M$ and hence $M^{g}=M$ by comparing orders. Since $N_{G}(M)=M$, we deduce that $g\in M$, contradicting the choice of $g.$
By Lemma 2.7, $G$ has a normal subgroup $K$ such that $G=MK$ and $M\cap K=H$. Since $M\neq G$ and $G$ is nonabelian simple, $K=G$ but then $M\cap K=M\neq H.$ This contradiction proves the theorem. ∎
Proof of Theorem 3..
In view of Lemma 2.11, it suffices to show that if $H$ is extremely closed in $G$, then $\langle H^{G}\rangle\cap N_{G}(H)=H$, where $H$ is a cyclic subgroup of $G$ of prime order $p>2.$ Suppose not and let $G$ be a counterexample to the claim with minimal order. Then $H=\langle x\rangle$ is a cyclic group of prime order $p>2$, $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G$ but $\langle H^{G}\rangle\cap N_{G}(H)\neq H.$
(1) We first claim that $\langle H^{G}\rangle=G.$ Suppose by contradiction that $L:=\langle H^{G}\rangle<G.$ By Lemma 2.2$(i),$ $H$ is pronormal in $G$ and hence $L=\langle H^{L}\rangle$ by Lemma 2.1$(vi).$ By Lemma 2.2$(ii),$ $H$ is extremely closed in $L$ and so by the minimality of $G,$ $\langle H^{L}\rangle\cap N_{L}(H)=H.$ However, as $L=\langle H^{L}\rangle,$ we have $N_{L}(H)=H,$ and thus
$\langle H^{G}\rangle\cap N_{G}(H)=L\cap N_{G}(H)=N_{L}(H)=H.$
This contradiction proves the claim.
(2) Let $N$ be a minimal normal subgroup of $G$. Then $N_{G}(H)N=HN=N_{G}(HN)$ and $N\not\leqslant H.$ Let $\overline{G}=G/N$ and use the ‘bar’ notation. By Lemma 2.2$(v),$ $\overline{H}$ is extremely closed in $\overline{G}$. By the minimality of $G,$ we have
$\langle\overline{H}^{\overline{G}}\rangle\cap N_{\overline{G}}(\overline{H})=\overline{H}$
and hence $N_{\bar{G}}(\bar{H})=\bar{H},$ as $\langle\bar{H}^{\bar{G}}\rangle=\bar{G}$ by (1). Thus $N_{G}(HN)=N_{G}(H)N=HN$ by appealing to Lemma 2.2$(iii).$ If $N\leqslant H$ then $N_{G}(H)=H$ and so $\langle H^{G}\rangle\cap N_{G}(H)=H,$ which is a contradiction. Therefore, $N\not\leqslant H$ as wanted.
(3) $O_{p}(G)=1$. Suppose not. Let $N\leqslant O_{p}(G)$ be a minimal normal subgroup of $G$. Then $N$ is an elementary abelian $p$-group and $HN$ is a $p$-subgroup of $G$ containing $H.$ By Lemma 2.2$(vi),$ $HN\leqslant N_{G}(H)$ and thus $N_{G}(H)=HN.$ As $N_{G}(H)$ contains a Sylow $p$-subgroup of $G$ by Lemma 2.2(vi), it follows that $HN\in{\mathrm{Syl}}_{p}(G)$ and $N_{G}(HN)=HN.$ Let $Q=H\cap N.$ Suppose that $Q>1.$ Since $N\leqslant N_{G}(H),$ by Lemma 2.3$(iii)$ we have $Q\unlhd G,$ forcing $Q=N$ by the minimality of $N.$ Hence $N\leqslant H,$ contradicting (2). Thus we can assume that $H\cap N=1.$ As $H,N$ are both normal in $HN,$ we have $[H,N]\leqslant H\cap N=1.$ Thus $HN$ is an abelian self-normalizing Sylow $p$-subgroup of $G.$ By Burnside’s normal $p$-complement theorem, $G=HNU,$ where $U=O_{p^{\prime}}(G).$ We have $\langle H^{G}\rangle=\langle H^{HU}\rangle\leqslant HU$. Now by (1), $G=\langle H^{G}\rangle=HU$ and thus $N=1,$ which is a contradiction.
(4) $O_{p^{\prime}}(G)=1$. Assume that $O_{p^{\prime}}(G)>1$ and let $N\leqslant O_{p^{\prime}}(G)$ be a minimal normal subgroup of $G$. Then $N\unlhd G$ is a $p^{\prime}$-group. Since $N_{G}(H)\leqslant HN$ and $N_{G}(H)$ contains a Sylow $p$-subgroup of $G,$ we deduce that $H\in{\mathrm{Syl}}_{p}(G).$ As $H\leqslant N_{G}(H)\leqslant HN,$ by Dedekind’s modular law, $N_{G}(H)=H(N_{G}(H)\cap N)=HK$ where $K=N_{N}(H)\unlhd N_{G}(H)$. Thus $[H,K]\leqslant H\cap K=1$, and so $N_{G}(H)=HK=C_{G}(H).$ Thus $H\leqslant Z(N_{G}(H)),$ and so by Burnside’s normal $p$-complement theorem, $G=HU,$ where $U$ is a normal $p^{\prime}$-subgroup of $G.$ By the Coprime group action theorem ([35, 8.2.7]), we have $U=[U,H]C_{U}(H)$ and $[U,H]=[U,H,H].$ As $H\cap U=1$ and $G=HU,$ we have $N_{G}(H)=C_{G}(H)=HC_{U}(H).$ Let $Q=[U,H].$ Then $G=HU=HC_{U}(H)Q$. As $G=\langle H^{G}\rangle=\langle H^{HC_{U}(H)Q}\rangle\leqslant HQ,$ we have $G=HQ$. By comparing orders, we obtain $U=Q=[U,H].$
Let $1\neq u\in U$ and let $T=\langle H,H^{u}\rangle.$ As $H$ is an abelian self-normalizing Sylow $p$-subgroup of $T$, we have $T=HO_{p^{\prime}}(T)$ and $C_{O_{p^{\prime}}(T)}(H)=1$. Moreover, $T=H[u,H],$ where $[u,H]=[T,x]\leqslant O_{p^{\prime}}(T)$ (for the first equality, see the remark after Theorem 2 in [14]). By comparing the orders, we get $[u,H]=O_{p^{\prime}}(T)$ and so $C_{[u,H]}(H)=1.$ By Lemma 2.9 (ii), we have
$C_{U}(H)=C_{[U,H]}(H)=\langle C_{[u,H]}(H):u\in[U,H]\rangle=1.$
Thus $N_{G}(H)=C_{G}(H)=HC_{U}(H)=H.$ Therefore, $\langle H^{G}\rangle\cap N_{G}(H)=H$, which is a contradiction.
Now let $N$ be a minimal normal subgroup of $G.$ Then $N\cong S^{k}$, where $k\geq 1$ is an integer and $S$ is a nonabelian simple group with $p$ dividing $|S|$.
[…]
(6) The final contradiction. Since $G=HN$ and $|H|=p$, either $H\cap N=1$ or $H\leqslant N$. The latter implies that $G$ is a finite nonabelian simple group.
(a) Assume $H\cap N=1$. Then $N_{G}(H)=C_{G}(H)$ or equivalently $N_{G}(\langle x\rangle)=C_{G}(x)$, where $H=\langle x\rangle$, Let $g\in G-C_{G}(x)$. Let $T=\langle H,H^{g}\rangle$. We will show that $T$ is a Frobenius group with Frobenius complement $H$. Indeed, since $H$ is extremely closed in $G$, we have
$T\cap N_{G}(H)=N_{T}(H)=H.$
Hence $H$ is a self-normalizing Sylow $p$-subgroup of $T$ of order $p.$ Thus $H\cap H^{t}=1$ for every $t\in T-H.$ So $T$ is a Frobenius group with complement $H$ as wanted. It follows that if $g\in G-C_{G}(x)$, then $x^{g}\neq x$ and $\langle x,x^{g}\rangle$ is a Frobenius group. Moreover, $G=\langle x^{G}\rangle$ by (1) and $N=G^{\prime}$ since $G=N\langle x\rangle$ and $N$ is perfect.
By Lemma 2.8, we have $C_{N}(x^{t})=1$ for some $t\in G,$ and then $C_{N}(x)=1$. Thus $C_{N}(H)=1$. We now have
$N_{G}(H)=C_{G}(H)=HC_{N}(H)=H.$
Therefore $\langle H^{G}\rangle\cap N_{G}(H)=H$, which is a contradiction.
(b) Assume that $G$ is a finite nonabelian simple group. We claim that $\langle x,x^{g}\rangle$ is solvable for all $g\in G.$ Let $g\in G$ and let $T=\langle H,H^{g}\rangle=\langle x,x^{g}\rangle.$ By Lemma 2.2$(iv),$ we have $T=HO_{p^{\prime}}(T).$ Since $N_{T}(H)=H,$ $H$ acts fixed point freely and coprimely on $O_{p^{\prime}}(T),$ the claim now follows from Lemma 2.10.
By Lemma 2.5, $p=3$ and $G$ is a finite simple group of Lie type defined over $\mathbb{F}_{3}$ or $G\cong\mathrm{U}_{n}(2)$ with $n\geq 4.$ Furthermore, except for $\mathrm{U}_{3}(3)$, the Lie rank of $G$ is strictly greater than $1$. Now it is easy to see that a Sylow $3$-subgroup of $G$ is nonabelian. Let $P$ be a Sylow $3$-subgroup of $G$ containing $x$. Then $H=\langle x\rangle$ is isolated in $P$ with respect to $G$, that is, $H$ does not conjugate in $G$ to any subgroup in $P-H$. By [22, Theorem 4.250], we deduce that $G\cong\mathrm{U}_{3}(3)$. By [25, Theorem A*], $x$ is a transvection. Now for any conjugate $x^{g}$ of $x$ different from $x$ and $x^{-1}$, we have that $T:=\langle x,x^{g}\rangle\cong{\mathrm{SL}}_{2}(3)$. However, $H$ is not self-normalizing in $T$. This final contradiction completes the proof. ∎
Proof of Theorem 4..
Let $H$ be an abelian $p$-subgroup of $G$ for some odd prime $p$. By Lemma 2.11, if $\langle H^{G}\rangle\cap N_{G}(H)=H$, then $G=N_{G}(H)O_{p^{\prime}}(G)$. Thus it remains to show that if $H$ is extremely closed in $G$, then $\langle H^{G}\rangle\cap N_{G}(H)=H$. Let $G$ be a counterexample to the claim above with $|G|$ minimal. Then $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G$ but $\langle H^{G}\rangle\cap N_{G}(H)\neq H.$ For the first two claims below, the arguments are similar to those in the proof of Theorem 3.
(1) $\langle H^{G}\rangle=G.$
(2) Let $N$ be a minimal normal subgroup of $G$. Then $N_{G}(H)N=HN=N_{G}(HN)$ and $N\not\leqslant H.$
(3) $F(G)=1.$ By way of contradiction, assume that $F(G)\neq 1.$ Then $G$ possesses a minimal normal subgroup $N$ which is an elementary abelian $r$-group for some prime $r.$ We now consider the following cases.
(a) $N$ is a $p$-group. We can use the same argument as in Claim (3) in the proof of Theorem 3 to obtain a contradiction.
(b) $N$ is a $p^{\prime}$-group. Using the same argument as in the first paragraph of the proof of Claim (4) in the previous theorem, we obtain that $H\in{\mathrm{Syl}}_{p}(G)$, $N_{G}(H)=C_{G}(H)$ and $G=HU,$ where $U$ is a normal $p$-complement of $G,$ and $U=[U,H].$
Recall that $\overline{G}=G/N$. As $N_{\overline{G}}(\overline{H})=\bar{H}$ and $\overline{G}=\overline{H}\overline{U},$ it follows that $C_{\overline{U}}(\overline{H})=1.$ By Lemma 2.10, $\overline{U}$ is solvable and so $U$ is solvable as $N$ is abelian. Let $1\neq u\in U$ and let $T=\langle H,H^{u}\rangle.$ Then $T=H[u,H],$ where $[u,H]=O_{p^{\prime}}(T)$ and $C_{[u,H]}(H)=1$ as $H$ is self-normalizing in $T.$ By Lemma 2.9(i), $C_{[U,H]}(H)=\langle C_{[u,H]}(H):u\in U\rangle=1.$ Thus $C_{U}(H)=C_{[U,H]}(H)=1$ and so $N_{G}(H)=C_{G}(H)=HC_{U}(H)=H.$ Therefore, $\langle H^{G}\rangle\cap N_{G}(H)=1$, which is a contradiction. Therefore, $F(G)=1$ as wanted.
Now let $N$ be a minimal normal subgroup of $G.$ Since $F(G)=1,$ $N$ must be isomorphic to a direct product of some non-abelian finite simple group $S.$
[…]
(5) $\langle H,H^{g}\rangle$ is solvable for all $g\in G.$ Let $g\in G$ and let $T=\langle H,H^{g}\rangle.$ By Lemma 2.2$(iv),$ we have $T=HO_{p^{\prime}}(T).$ Since $N_{T}(H)=H,$ $H$ acts fixed point freely and coprimely on $O_{p^{\prime}}(T),$ the claim now follows from Lemma 2.10.
(6) $G$ is an almost simple group with simple socle $S$ and $G=SH$. Since $G=HN$, where $N\cong S^{k}$ for some integer $k\geq 1$ and $S$ is a nonabelian simple group. It suffices to show that $N$ is simple.
Let $W\unlhd N$ be a component of $G.$ Assume that $W\neq N.$ As $W\unlhd N,$ $N\leqslant N_{G}(W)$ and so $N_{G}(W)=N_{H}(W)N.$ Since $W<N,$ $N_{G}(W)<G$ and so $N_{H}(W)<H.$ Let $x\in H-N_{H}(W).$ Then $x\not\in N_{G}(W)$ and $x^{2}\not\in C_{G}(W)$ since $p$ is odd, whence $x$ and $W$ satisfy the hypothesis of Lemma 2.4 and hence $\langle x,x^{g}\rangle$ is not solvable for some $g\in G$ but then this contradicts (5) as $\langle x,x^{g}\rangle\leqslant\langle H,H^{g}\rangle$ is solvable. Thus $W\unlhd G$ and so $G=HN,$ where $N$ is a nonabelian simple group which is also a minimal normal subgroup of $G.$ Thus $G$ is an almost simple group with socle $S$ as wanted.
(7 ) The final contradiction.
By (5) and (6), $G=HS$ is an almost simple group with socle $S$ and if $x\in H$ is an element of order $p$, then $\langle x,x^{g}\rangle$ is solvable for all $g\in G$. By Lemma 2.5, $p=3$ and $S$ is a finite simple group of Lie type defined over $\mathbb{F}_{3}$ or $S\cong\mathrm{U}_{n}(2),n\geqslant 4.$ By Theorem 3, we may assume that $|H|\geqslant 9.$ Since $G=HS$, we have $G/S\cong H/(H\cap S)$ is a $3$-group and thus if $G\neq S$, then elements in $H-S$ induce outer automorphisms of $S$ of $3$-power order.
$(i)$ Assume $S\cong\mathrm{U}_{n}(2),n\geqslant 4$ or $G\neq S.$ If $S\cong\mathrm{U}_{n}(2),$ then since $n\geqslant 4,$ we have $|\operatorname{Out}(\mathrm{U}_{n}(2))|=2(n,3)$. If $G\neq S$ and $S\not\cong\mathrm{U}_{n}(2)$, then $S\cong{\rm D}_{4}(3)$ or ${}^{3}{\rm D}_{4}(3)$ since $S$ has no nontrivial field automorphism so $\operatorname{Out}(S)$ contains diagonal automorphisms and possibly graph automorphisms only. In all cases, the Sylow $3$-subgroup of $\operatorname{Out}(S)$ has order at most $3$. Hence $|G:S|=|H:(S\cap H)|\leqslant 3$.
Since $|H|\geq 9$, if $G$ is not simple then $A=H\cap S>1.$ By Lemma 2.3(ii), $A$ is strongly closed in $S.$ If $G=S=U_{n}(2)$, then $A=H$ is strongly closed in $S$. In both cases, $S$ contains a nontrivial abelian $3$-subgroup $A$ which is strongly closed in $S.$ Since $S$ is simple, $S=\langle A^{S}\rangle.$ As a Sylow $3$-subgroup of $S$ is nonabelian, $A$ cannot be a Sylow $3$-subgroup of $S$; however this contradicts [15, Theorem 1.2 (i)].
$(ii)$ $G=S$ is a finite simple group of Lie type defined over $\mathbb{F}_{3}$. The possibilities for $G$ can be read off from Table 1 in [25] and note the correction in Remark 5.2 in [37].
Let $P$ be a Sylow $p$-subgroup of $G$ containing $H,$ and let $M$ be a maximal parabolic subgroup of $G$ containing the Borel subgroup $B:=N_{G}(P)$. Let $R=O_{p}(M).$ Then $M=N_{G}(R)$. By induction, we have $\langle H^{M}\rangle\cap N_{M}(H)=H,$ and hence $\langle H^{M}\rangle=HU,$ where $U=O_{p^{\prime}}\langle H^{M}\rangle$ and $C_{U}(H)=1.$ Therefore, $U$ is a solvable normal $p^{\prime}$-subgroup of $M$ since $\langle H^{M}\rangle\unlhd M$. As $H\leqslant\langle H^{M}\rangle\unlhd M$ and $H$ is pronormal in $M$, we have $M=N_{M}(H)U$. However, as $M$ is a parabolic subgroup of a finite simple group of Lie type $G$, $F^{*}(M)=O_{p}(M)$ (Corollaries 3.1.4 and 3.1.5 in [23]), so $U=1.$ Hence $H\unlhd M$ and $M=N_{G}(H)$. Thus $M$ contains every maximal parabolic subgroup of $G$ that contains the Borel subgroup $B.$ However this happens only when the Lie rank of $G$ is $1$ Therefore $G\cong\mathrm{U}_{3}(3).$ (We can also use Theorem 1.2 in [15] to arrive at this conclusion.)
From the Atlas [7], we have $P\cong 3_{+}^{1+2}$ is an extraspecial group of order $27$ and exponent $3$. Thus $|H|=9$. Since $G$ has Lie rank $1$, the Borel subgroup $B$ of $G$ is a maximal subgroup of $G$. Hence $B=N_{G}(H).$ Let $g\in G-B$ and $T=\langle H,H^{g}\rangle$. We know that $T=HO_{3^{\prime}}(T)$ is solvable, so $T<G$ and thus $T$ lies in some maximal subgroup of $G$ whose order must be divisible by $|H|=9.$ Inspecting the list of maximal subgroups of $G$ in the Atlas [7], the only maximal overgroups of $H$ and $T$ in $G$ are the Borel subgroups $B$ and its $G$-conjugates. Hence $\langle H,H^{g}\rangle\leqslant B^{t}$ for some $t\in G.$ Note that $B=P:W$, where $W=\langle y\rangle$ is a cyclic group of order $8$. Since $B^{t}$ has a normal Sylow $3$-subgroup, we must have that $H\leqslant P^{t}\unlhd B^{t}$ and hence $H$ is subnormal in $B^{t}.$ Since $H$ is pronormal in $G$ and hence in $B^{t},$ we have $H\unlhd B^{t}$ or $B^{t}=N_{G}(H)=B$. Therefore, $\langle H,H^{g}\rangle\leqslant B$ for all $g\in G.$ However, this implies that $\langle H^{G}\rangle\leqslant B$, which is a contradiction. The proof is now complete. ∎
Proof of Theorem 5..
Let $G$ be a counterexample to Theorem 5 with $|G|$ minimal. Then $\langle H,H^{g}\rangle\cap N_{G}(H)=H$ for all $g\in G$ but $G\neq N_{G}(H)O_{2^{\prime}}(G)$, where $H$ is an abelian $2$-subgroup of $G$.
(1) $\langle H^{G}\rangle=G.$ Let $L=\langle H^{G}\rangle.$ Assume $L\neq G.$ By Lemmas 2.2$(i)$ and 2.1$(iii),$ $G=N_{G}(H)L.$ By Lemma 2.2$(ii)$ and the minimality of $G,$ we have $L=N_{L}(H)O_{2^{\prime}}(L).$ As $L\unlhd G,$ $O_{2^{\prime}}(L)\leqslant O_{2^{\prime}}(G),$ and hence
$G=N_{G}(H)L=N_{G}(H)O_{2^{\prime}}(L)=N_{G}(H)O_{2^{\prime}}(G).$
This contradictions shows that $L=G.$
(2) $O_{2^{\prime}}(G)=1.$ Suppose by contradiction that $N=O_{2^{\prime}}(G)\neq 1.$ It follows from Lemma 2.2$(v)$ that the hypothesis carries over to $\overline{G}=G/N,$ and so by the minimality of $G,$ we obtain $\overline{G}=N_{\overline{G}}(\overline{H})O_{2^{\prime}}(\overline{G}).$ But $O_{2^{\prime}}(\overline{G})=O_{2^{\prime}}(G/O_{2^{\prime}}(G))=1,$ and hence $\overline{G}=\overline{N_{G}(H)},$ by Lemma 2.2$(iii).$ Hence $G=N_{G}(H)N=N_{G}(H)O_{2^{\prime}}(G),$ which is a contradiction. This proves the claim.
(3) $O_{2}(G)=1.$ Assume $O_{2}(G)\neq 1.$ Let $N\leqslant O_{2}(G)$ be a minimal normal subgroup of $G$ and let $U=H\cap N.$ Observe first that $O_{2}(G)\leqslant N_{G}(H)$ since $P\leqslant N_{G}(H)$ for any $H\leqslant P\in{\mathrm{Syl}}_{2}(G).$ It follows that $N\leqslant N_{G}(H).$ We will show that $N\leqslant Z(G).$ Suppose first that $U\neq 1.$ Lemma 2.3(iii) implies that $U\unlhd G.$ By the minimality of $N,$ we have $U=N,$ and so $N\leqslant H.$ As $H$ is abelian, $H\leqslant C_{G}(N)\unlhd G.$ By (1), we have $C_{G}(N)=G,$ and $N\leqslant Z(G).$ If $U=H\cap N=1,$ then as $H$ and $N$ are both normal in $N_{G}(H),$ we have $[H,N]\leqslant H\cap N=1$ and thus $H\leqslant C_{G}(N)\unlhd G,$ so $N\leqslant Z(G).$
[…]
$G=N_{G}(H)K=N_{G}(H)N=N_{G}(H).$
This contradiction proves the claim.
(4) $G$ is finite nonabelian simple group. Let $N$ be a minimal normal subgroup of $G.$ By (2) and (3), $N\cong S^{k}$ for some finite nonabelian simple group $S$ and some integer $k\geq 1$. Let $L=HN.$ Suppose that $L<G.$ By the minimality of $G$, we have $L=N_{L}(H)O_{2^{\prime}}(L).$ Hence $O_{2^{\prime}}(L)\leqslant N,$ and so $O_{2^{\prime}}(L)=1,$ since $N$ is a product of non-abelian simple groups. We deduce that $H\unlhd L.$ As $H$ is an abelian $2$-group and $N\cong S^{k}$, we have $H\cap N=1$ and hence $[H,N]\leqslant H\cap N=1$. Thus $H\leqslant C_{G}(N)\unlhd G.$ By (1), we have $C_{G}(N)=G,$ and then $N\leqslant Z(G).$ This contradiction shows that $G=HN.$ Thus $F^{*}(G)=N$ and $H$ is a strongly closed abelian $2$-subgroup of $G$. Now by [20, Theorem A], we have $F^{*}(G)=G$ and so $N=G.$ It follows that $G=S$ is simple.
(5) The final contradiction. By [20, Theorem A] again, $G$ is isomorphic to one of the following groups:
$(1)$ $\mathrm{L}_{2}(2^{n}),n\geqslant 3;$ ${}^{2}{\rm B}_{2}(2^{2n+1}),n\geqslant 1;$ or $\mathrm{U}_{3}(2^{n}),n\geqslant 2.$
[…]
$(3)$ ${}^{2}{\rm G}_{2}(3^{2n+1}),n\geqslant 1;$ or ${\rm J}_{1},$ the first Janko group.
By Glauberman’s $Z^{*}$-theorem, we may assume that $|H|\geq 4.$
(a) Assume $G$ is isomorphic to one of the groups in $(1).$ Let $H\leqslant P\in{\mathrm{Syl}}_{2}(G)$ and let $B=PT$ be the Borel subgroup of $G$ containing $P.$ By [20, (3.2)], $H=Z(P)$ is a non-cyclic elementary abelian $2$-group, $P$ is a $T.I$ subgroup of $G,$ and $B$ is the unique maximal subgroup of $G$ containing $P.$ It follows that $B=N_{G}(H)$. Observe that for any $1\neq x\in H,$ $P\leqslant C_{G}(x)\leqslant B,$ as $P$ is uniquely contained in $B.$ For $g\in G-B,$ let $T=\langle H,H^{g}\rangle.$ Then $N_{T}(H)=H$ and then by Burnside’s normal $p$-complement theorem, we have $T=HU,$ where $U=O_{2^{\prime}}(T)\unlhd T.$ As $H$ is non-cyclic abelian, $U=\langle C_{U}(a):1\neq a\in H\rangle$ (see [35, 8.3.4]). However, as $C_{G}(a)\leqslant B$ for all $1\neq a\in H,$ we have $U\leqslant B$ and so $T=\langle H,H^{g}\rangle\leqslant B.$ From the hypothesis, we must have $T=T\cap B=H,$ hence $H^{g}=H.$ This implies that $g\in B,$ contradicting the choice of $g.$
(b) Assume that $G\cong\mathrm{L}_{2}(5)\cong{\mathrm{A}}_{5}.$ By [20, (3.4)], $H\in{\mathrm{Syl}}_{2}(G),$ $N_{G}(H)\cong{\mathrm{A}}_{4}$ and $N_{G}(H)$ is the unique maximal subgroup of $G$ containing $C_{G}(a),$ for all $1\neq a\in H.$ Take $g\in G-N_{G}(H),$ and let $T=\langle H,H^{g}\rangle.$ Then $T=HU,$ where $U=O_{2^{\prime}}(T).$ As $H$ is non-cyclic, $U=\langle C_{U}(a):1\neq a\in H\rangle\leqslant N_{G}(H).$ This leads to a contradiction as in the previous case.
(c) Assume $G$ is isomorphic to one of the groups in $(2)$ with $q\geqslant 11.$ By [20, (3.4)], we have $H\in{\mathrm{Syl}}_{2}(G)$ and $N_{G}(H)\cong{\mathrm{A}}_{4}.$ Clearly, $G$ contains a maximal subgroup $M$ isomorphic to the dihedral group $D_{q\pm 1}$ such that $M$ does not contain $H$. Assume $M$ is generated by two involutions $a,b.$ We can choose $a,b$ such that $b\in H.$ Now $G=\langle M,H\rangle\leqslant\langle a,H\rangle,$ and hence $G=\langle a,H\rangle.$ By Sylow’s Theorem, there exists some $g\in G$ such that $a\in H^{g}.$ Thus $G=\langle a,H\rangle\leqslant\langle H^{g},H\rangle,$ and then $G=\langle H,H^{g}\rangle,$ which contradicts the hypothesis that $\langle H,H^{g}\rangle\cap N_{G}(H)=H.$
(d) Finally, assume that $G$ is isomorphic to one of the groups in $(3).$ By [20, (3.4)], we have $H\in{\mathrm{Syl}}_{2}(G).$ Now $H$ contains an involution $t$ such that $C_{G}(t)=\langle t\rangle\times L,$ where $L\cong\mathrm{L}_{2}(q),$ $q\equiv 3,5$ (mod $8$) and $[G:C_{G}(t)]$ is odd (see [41]). As $H\leqslant C_{G}(t)<G,$ by the minimality of $G$, we obtain $C_{G}(t)=(C_{G}(t)\cap N_{G}(H))O_{2^{\prime}}(C_{G}(t)).$ However, as $C_{G}(t)=\langle t\rangle\times L,$ where $L$ is nonabelian simple, it follows that $O_{2^{\prime}}(C_{G}(t))=1$ whence $C_{G}(t)\leqslant N_{G}(H)$. Hence $C_{G}(t)=N_{G}(H)$ since $C_{G}(t)$ is maximal in $G$ by [20, $(3.4)$]. It follows that $H\unlhd C_{G}(t)$ and then $H\cap L\unlhd L,$ where $H\cap L\in{\mathrm{Syl}}_{2}(L),$ which contradicts the simplicity of $L.$ This final contradiction proves the theorem. ∎
Remark 3.1.
Cases (b)-(d) in the proof of Theorem 5 can be eliminated using a result due to Guralnick in [26] stating that every finite nonabelian simple group can be generated by an involution and a Sylow $2$-subgroup.
Proof of Corollary 1.
Let $H$ be an extremely closed abelian $p$-subgroup of $G$ for some prime $p$. Assume first that $p=2$. By Theorem 5, $G=N_{G}(H)O_{2^{\prime}}(G)$ which implies that $HO_{2^{\prime}}(G)\unlhd G$ and clearly $HO_{2^{\prime}}(G)$ is solvable and thus $H\subseteq R(G).$ Assume now that $p$ is odd. By Theorem 4, we have $N_{G}(H)\cap\langle H^{G}\rangle=H$ which implies that $H$ is a self-normalizing abelian Sylow $p$-subgroup of $L:=\langle H^{G}\rangle$. By Burnside’s normal $p$-complement theorem, $L=HO_{p^{\prime}}(L)$ and $C_{O_{p^{\prime}}(L)}(H)=1.$ By Lemma 2.10, $O_{p^{\prime}}(L)$ is solvable and thus $L=HO_{p^{\prime}}(L)$ is a solvable normal subgroup of $G$. This implies that $H\leqslant R(G)$ as wanted. The proof is now complete. ∎
4. A $p$-solvability criterion
Let $p$ be a prime. A finite group $G$ is said to be a minimal non-$p$-solvable group if $G$ is not $p$-solvable but every proper subgroup of $G$ is $p$-solvable. A minimal simple group is a nonabelian finite simple groups whose all proper subgroups are solvable. Observe that minimal non-$2$-solvable simple groups are exactly the minimal simple groups and these groups are classified by Thompson in [39].
Lemma 4.1.
Every minimal simple group is isomorphic to one of the following simple groups:
  • $(1)$
    $\mathrm{L}_{2}(2^{r}),r$ is a prime.
  • $(2)$
    $\mathrm{L}_{2}(3^{r}),r$ is an odd prime.
  • $(3)$
    $\mathrm{L}_{2}(r),r>3$ is a prime such that $15\mid r^{2}+$ 1.
  • $(4)$
    ${}^{2}{\rm B}_{2}(2^{r}),r$ is an odd prime.
  • $(5)$
    $\mathrm{L}_{3}(3).$
Proof.
This is [39, Corollary 1]. ∎
The next result classifies minimal non-$3$-solvable simple groups.
Lemma 4.2.
Let $G$ be a finite non-abelian simple group. Assume that every proper subgroup of $G$ is $3$-solvable. Then $G$ is isomorphic to a minimal simple group or the Suzuki group ${}^{2}{\rm B}_{2}(q)$ with $q=2^{2m+1},m\geq 1.$
Proof.
This is Lemma 5.3 in [29]. ∎
Finally, we need the classification of finite non-$p$-solvable simple groups for any primes $p\geq 5$.
Lemma 4.3.
Let $G$ be a finite nonabelian simple group and let $p\geqslant 5$ be a prime dividing $|G|$. Assume that every proper subgroup of $G$ is $p$-solvable. Then one of the following holds.
  • $(1)$
    $G=\mathrm{L}_{2}(p)$.
  • $(2)$
    $G={\mathrm{A}}_{p}.$
  • $(3)$
    $G=\mathrm{L}_{2}(q)$ with $p\mid q^{2}-1$.
  • $(4)$
    $G=\mathrm{L}_{n}(q),n\geq 3$ is odd, and $p$ divides $q^{n}-1$ but not $\prod_{i=1}^{n-1}(q^{i}-1)$.
  • $(5)$
    $G=\mathrm{U}_{n}(q),n\geq 3$ is odd, and $p$ divides $q^{n}-(-1)^{n}$ but not $\prod_{i=1}^{n-1}(q^{i}-(-1)^{i})$.
  • $(6)$
    $G={}^{2}{\rm B}_{2}(q)$ with $q=2^{2m+1},m\geq 1.$
  • $(7)$
    $G={}^{2}{\rm G}_{2}(q)$ and $p\mid(q^{2}-q+1)$, where $q=3^{2m+1},m\geq 1.$
  • $(8)$
    $G={}^{2}{\rm F}_{4}(q)$ with $q=2^{2m+1},m\geq 1$ and $p\mid(q^{4}-q^{2}+1).$
  • $(9)$
    $G={}^{3}{\rm D}_{4}(q)$ and $p\mid(q^{4}-q^{2}+1).$
  • $(10)$
    $G={\rm E}_{8}(q)$ and $p$ divides $(q^{30}-1)$ but not $\prod_{i\in\{8,14,18,20,24\}}(q^{i}-1)$.
  • $(11)$
    $(G,p)$ is one of the following $({\rm M}_{23},23)$, $({\rm J}_{1},7\text{ or }19)$, (Ly,37 or $({\rm Ly},37\text{ or }67)$, $({\rm J}_{4},29\text{ or }43)$, $({\rm Fi}_{24}^{\prime},29)$, $({\rm B},47)$ or (M,41 or 59 or $({\rm M},41\text{ or }59\text{ or }71).$
Proof.
This is Lemma 5.4 in [29]. ∎
Let $q$ be a prime power and let $n\geq 2$ be an integer. A prime divisor $p$ of $q^{n}-1$ is called a primitive prime divisor or ppd of $q^{n}-1$ if $p$ does not divide $q^{k}-1$ for all integers $k$ with $1\leqslant k<n.$ Zsigmondy’s theorem [43] states that such a ppd p exists unless $(n,q)=(6,2)$ or $n=2$ and $q$ is a Mersenne prime. Now if $n>1$ is an integer and $p$ is a prime, then the $p$-part of $n,$ denoted by $n_{p}$, is the largest power of $p$ dividing $n.$ We refer the reader to [34, 5] for the description of maximal subgroups of finite simple groups of Lie type.
Proposition 4.4.
Let $G$ be a finite nonabelian simple group and $p$ be a prime dividing $|G|$. Assume that every proper subgroup of $G$ is $p$-solvable. Let $P$ be a Sylow $p$-subgroup of $G$. Then $G=\langle P,P^{g}\rangle$ for some $g\in G.$
Proof.
Let $G$ be a finite nonabelian simple group and let $p$ be a prime dividing $|G|$. Let $P\in{\mathrm{Syl}}_{p}(G)$. Assume that every proper subgroup of $G$ is $p$-solvable. Let $x\in P$ with $|x|=p$. Then $\langle x,x^{g}\rangle\leqslant\langle P,P^{g}\rangle$ for all $g\in G$ and thus $\langle x,x^{g}\rangle$ is $p$-solvable for all $g\in G.$
(a) If $p=2$, then every finite nonabelian simple group $G$ can be generated by two Sylow $2$-subgroups by Theorem A in [26]. So, we may assume that $p>2$.
(b) If $G$ is a finite nonabelian simple group of Lie type in characteristic $p$, then $G$ is generated by two Sylow $p$-subgroups by Proposition 2.5 in [6].
(c) Assume that $p=3$. By Lemma 4.2, since $3$ divides $|G|$, $G$ is a minimal simple group. By part (b), we only need to consider the cases when $G$ is isomorphic to $\mathrm{L}_{2}(2^{r}),r$ is a prime, or $\mathrm{L}_{2}(r),r>3$ is a prime and $15\mid r^{2}+$ 1.
If $G\cong\mathrm{L}_{2}(4)$, then we can check by using GAP [17] that there exists $g\in G$ such that $G=\langle P,P^{g}\rangle$. Assume next that $G\cong\mathrm{L}_{2}(q)$, $q=2^{r}$ or $r,$ where $r$ is an odd prime. By [25, Theorem A*], there exists an element $x\in G$ of order $3$ such that $\langle x,x^{g}\rangle$ is nonsolvable for some $g\in G.$ Since $G$ is minimal simple, we must have $G=\langle x,x^{g}\rangle$ and thus $G=\langle P,P^{g}\rangle.$
(d) Assume that $p\geq 5.$ By part (b), and Lemma 4.3, $G$ is one of the groups listed in $(2)-(11)$ in that lemma. We now consider each case in turn.
(1) Assume $G={\mathrm{A}}_{p}$. In this case $|P|=p$. Without loss of generality, take $P=\langle x\rangle$, where $x=(1,2,\dots,p)$ is a $p$-cycle in ${\mathrm{A}}_{p}.$ Let $y=(1,2,p,p-1,p-2,\dots,3)\in{\mathrm{A}}_{p}$ be another $p$-cycle. Then $xy=(1,3,2)$ and clearly $\langle x,y\rangle=\langle(xy)^{-1},x\rangle=\langle(1,2,3),(1,2,\dots,p)\rangle={\mathrm{A}}_{p}$ (see, for example, [33, Theorem B].) Hence $G$ is generated by two Sylow $p$-subgroups.
(2) Assume $G=\mathrm{L}_{2}(q)$ with $p\mid q^{2}-1$. If $q\leqslant 11$, we can check using GAP [17] that the result holds. Assume $q\geq 13.$ Inspecting the argument in [25, §5.1.2], if $x$ is any element of order $p$, then we can find $g\in G$ such that $\langle x,x^{g}\rangle\cong\mathrm{L}_{2}(q)$ and hence $\langle P,P^{g}\rangle=G$.
(3) Assume $G=\mathrm{L}_{n}(q),n\geq 3$ is odd, and $p$ divides $q^{n}-1$ but not $\prod_{i=1}^{n-1}(q^{i}-1)$. Write $q=s^{f}$, where $s$ is a prime and $f\geq 1$ is an integer. In this case $p$ is a ppd of $q^{n}-1$. Hence $P\in{\mathrm{Syl}}_{p}(G)$ is cyclic of order $(q^{n}-1)_{p}.$ Since $n\equiv 1$ mod $p$ and $n\geq 3$ is odd, we have $p\geq 2n+1$ and $p\nmid n$.
Assume that $t$ is a prime divisor of $n$ and write $n=tm$ for some integer $m\geq 1.$ Assume that $m>1$. Then $G$ has a $\mathcal{C}_{3}$-subgroup $H$ of type ${\mathrm{GL}}_{m}(q^{t})$ (see [34, Table 3.5A]) which is maximal and contains a Sylow $p$-subgroup of $G$. Since $n\geq 3$ is odd, $m,t\geq 3$ and so $q^{t}\geq 2^{t}\geq 8$. Therefore, $H$ is not $p$-solvable.
Thus we can assume that $n=t$ is an odd prime. If $P$ lies in a unique maximal subgroup $H$ of $G$, then $H$ is of type ${\mathrm{GL}}_{1}(q^{n})$ by [2, Table B]. We can choose $g\in G-H$ such that $P^{g}\not\leqslant H$ and hence $G=\langle P,P^{g}\rangle$. Assume that $P$ lies in some other maximal subgroup $M$ of $G$ not of type ${\mathrm{GL}}_{1}(q^{n})$. As in the proof of Case 3 of Proposition 6.2 in [2], $M\in\mathcal{C}_{5}$ is a subfield subgroup of type ${\mathrm{GL}}_{n}(q_{0})$, where $q=q_{0}^{k}$, $k$ is an odd prime and $(q_{0}^{n}-1)_{p}=(q^{n}-1)_{p}$ or $M\in\mathcal{S}$ is almost simple with socle $S\cong\mathrm{L}_{2}(p)$ and $n=(p-1)/2$. However, in both cases, $M$ is not $p$-solvable.
(4) Assume $G=\mathrm{U}_{n}(q),n\geq 3$ is odd, and $p$ divides $q^{n}+1$ but not $\prod_{i=1}^{n-1}(q^{i}-(-1)^{i})$. Write $q=s^{f}$ where $s$ is a prime and $f\geq 1$.
(a) Assume that $n=3$. If $q=3$, then $p=7$. In this case, $|P|=7$ and $P$ lies in $\mathrm{L}_{2}(7)$ which is not $7$-solvable. Similarly, if $q=5$, then $p=7$ and $|P|=7$ and $P$ lies in ${\mathrm{A}}_{7}$.
Assume first that $q$ is a prime. Let $H$ be a maximal subgroup of $G$ containing $P.$ By the proof of Proposition 6.3 in [2], either $P$ lies in a uniqiue maximal subgroup of $G$ and we are done or $P$ is contained in $\mathrm{L}_{2}(7)$ and $p=7$; however, $\mathrm{L}_{2}(7)$ is not $7$-solvable.
Assume $q=s^{f}$ with $f>1$. In this case, if $P$ is not contained in a unique maximal subgroup, then $P$ can be contained in a subfield subgroup of type $\mathrm{GU}_{3}(q_{0})$ with $q=q_{0}^{k}$, and $k$ is an odd prime (see [5, Table 8.5]). However such a maximal subgroup is not $p$-solvable.
(b) Assume $n\geq 5.$ Then $p$ is a ppd of $q^{2n}-1$. Hence $p\geq 2n+1$.
Assume that $n=tm$, where $t$ is a prime divisor of $n$ and $m>1$. Since $n\geq 5$ is odd, $t,m\geq 3.$ Then $G$ has a maximal subgroup of type $\mathrm{GU}_{m}(q^{t})$ and contains a Sylow $p$-subgroup of $G$. Since $q^{t}\geq 2^{t}\geq 8$, such a maximal subgroup is not $p$-solvable.
Therefore, $n=t\geq 5$ is a prime. Argue as in case (3), if $P$ lies in a unique maximal subgroup of $G$, then the conclusion holds. As in the proof of Proposition 6.4 [2], $P$ lies in a subfield subgroup $H$ of type $\mathrm{GU}_{n}(q_{0})$, where $q_{0}^{k}=q$ and $k\geq 3$ is a prime, or of type $\mathrm{O}_{n}(q)$ or $H$ is an almost simple group with socle $\mathrm{L}_{2}(p)$ with $n=(p-1)/2$ and $7\leqslant p\equiv 3$ mod 4. However, in all cases, these maximal subgroups are not $p$-solvable.
(5) $G={}^{2}{\rm B}_{2}(q)$ with $q=2^{2m+1},m\geq 1.$ Then $|G|=q^{2}(q-1)(q+s+1)(q-s+1),$ where $s=\sqrt{2q}=2^{m+1}$. The maximal subgroups of $G$ are listed in [5, Table 8.16]. Since $p$ is not the characteristic of $G$, $p>2$ and $p\mid q-1$ or $p\mid q\pm s+1$.
Assume first that $p\mid q-1$. Then $P$ lies in maximal subgroups of the form $[q^{2}]:(q-1)$ and $D_{2(q-1)}$. It follows that $\langle P,p^{g}\rangle$ is solvable for all $g\in G.$ Let $x\in P$ with $|x|=p\geq 5.$ Then $\langle x,x^{g}\rangle$ is solvable for all $g\in G$. However, this is impossible in view of Theorem A* in [25].
Assume that $p\mid q\pm s+1.$ In this case, $P$ lies in a maximal subgroup of the form $(q\pm s+1):4$ or a subfield subgroup of the form ${}^{2}{\rm B}_{2}(q_{0})$, where $q_{0}^{k}=q,k\geq 3$ is a prime and $q_{0}>2.$ Clearly, the subfield subgroup is not $p$-solvable (if it contains $P$). Hence $P$ lies in a unique maximal subgroup of $G$ and the result follows.
(6) $G={}^{2}{\rm G}_{2}(q)$ with $q=3^{2m+1},m\geq 1$ and $p\mid q^{2}-q+1$. We can use the same argument as in the previous case using [5, Table 8.43].
(7) $G={}^{2}{\rm F}_{2}(q)$ with $q=2^{2m+1},m\geq 1$ and $p\mid q^{4}-q^{2}+1$. In this case, $p$ is a ppd of $q^{12}-1$. Using the argument in Proposition 7.2 in [2], either $P$ lies in a unique maximal subgroup or it lies in a subfield subgroup ${}^{2}{\rm F}_{2}(q_{0})$, which is not $p$-solvable.
(8) $G={}^{3}{\rm D}_{4}(q)$ and $p\mid q^{4}-q^{2}+1$. We can use the argument in Proposition 7.3 in [2] to obtain the conclusion as in the previous case.
(9) $G={\rm E}_{8}(q)$ and $p$ divides $(q^{30}-1)$ but not $\prod_{i\in\{8,14,18,20,24\}}(q^{i}-1)$. In this case, $p$ is a ppd of $q^{30}-1$. From Proposition 7.10 [2], either $P$ lies in a unique maximal subgroup and the result follows or $P$ can lie in a maximal exotic local subgroup $2^{5+10}\cdot\mathrm{L}_{5}(2)$ when $|P|=p=31$ or $P$ lies in an almost simple group. In the last two possibilities, clearly, these maximal subgroups are not $p$-solvable.
(10) $(G,p)$ is one of the following $({\rm M}_{23},23)$, $({\rm J}_{1},7\text{ or }19)$, (Ly,37 or $({\rm Ly},37\text{ or }67)$, $({\rm J}_{4},29\text{ or }43)$, $({\rm Fi}_{24}^{\prime},29)$, $({\rm B},47)$ or (M,41 or 59 or $({\rm M},41\text{ or }59\text{ or }71).$
By [2, Table D], $P$ lies in the unique maximal subgroup of $G$ and the result follows except for the case $(G,p)=({\rm J}_{1},7)$. By the Atlas [7], the maximal subgroups of ${\rm J}_{1}$ containing a Sylow $7$-subgroup are isomorphic to either $32^{3}:7:$ 3 or $7:6$. Thus $\langle x,x^{g}\rangle$ is solvable for all $g\in{\rm J}_{1}$, where $x\in P$ with $|x|=7$. However, this contradicts Theorem A* in [25]. The proof is now complete. ∎
Remark 4.5.
It is conjectured in [6] that if $G$ is a finite nonabelian simple group and if $r$ and $s$ are prime divisors of $|G|$, then $G$ can be generated by a Sylow $r$-subgroup and a Sylow $s$-subgroup. The previous proposition is just a special case of this conjecture when $r=s=p$ and $G$ is a minimal non-$p$-solvable finite simple group.
Proof of Theorem 6.
Let $G$ be a finite group and let $p$ be a prime. Let $P$ be a Sylow $p$-subgroup of $G$. If $G$ is $p$-solvable, then every subgroup of $G$ is $p$-solvable. Therefore, it suffices to show that if $\langle P,P^{g}\rangle$ is $p$-solvable for all $g\in G,$ then $G$ is $p$-solvable. Suppose not and let $G$ be a counterexample with $|G|$ minimal. Then $\langle P,P^{g}\rangle$ is $p$-solvable for all $g\in G$ but $G$ is not solvable.
We first claim that every proper subgroup of $G$ is $p$-solvable and thus $G$ is a minimal non-$p$-solvable group. Let $H$ be a proper subgroup of $G$ and let $Q$ be a Sylow $p$-subgroup of $H$. Then $Q\leqslant P^{t}$ for some $t\in G.$ Now for every $h\in H,$ we have
$\langle Q,Q^{h}\rangle\leqslant\langle P^{t},(P^{t})^{h}\rangle=\langle P,P^{tht^{-1}}\rangle^{t}.$
Since $\langle P,P^{tht^{-1}}\rangle$ is $p$-solvable, $\langle Q,Q^{h}\rangle$ is $p$-solvable. Therefore, by the minimality of $|G|$, $H$ is $p$-solvable.
By Proposition 4.4, we know that $G$ is not a nonabelian simple group. Let $N$ be a proper nontrivial normal subgroup of $G$. Now $PN/N$ is a Sylow $p$-subgroup of $G/N$ and it satisfies the hypothesis of the theorem. Since $|G/N|<|G|$, $G/N$ is $p$-solvable. As in the previous claim, $N$ is also $p$-solvable and thus $G$ is $p$-solvable as well. This final contradiction proves the theorem. ∎